If separate samples of argon, methane, nitrogen and carbon dioxide, all at the same initial temperature and pressure are expanded adiabatically reversibly to double their original volume, then which one of these gases will have high temperature as final temperature?

Grade:11

1 Answers

Susmita
5 years ago
Initially they are at the same temperature.Finally V2/V1=2 for all of them.
PVgamma=constant
Or,TVgamma-1=constant
So
T2=T1(V1/V2)gamma-1=T1(1/2)gamma-1
T1 is same for all.T2 will be highest for that one whose gamma is least.gamma is least for CO2 here.
If you want to check yourself do it with the formula
Gamma=1+(2/f)
Where f is the degree of freedom and
f=3N-m,where N is number of atoms in a molecule and m is number of constraints.
If it helps please approve.